### v2

@@ Appendix H Centaur Experiments
    Self-persuasion [8]. When prompted to consider the other side, do digital twins abandon their attitudes more readily than their human counterparts?
@@
In sum, across all three task types, we do not find evidence supporting the possibility that because twins obediently follow orders, they are more sensitive to survey instructions, such as "consider the alternative point of view" or "imagine…" If anything, we find that humans, on average, respond more to instructions in this survey than digital twins.
@@
Building on the seminal work of [1], this study assesses the ability of digital twins to replicate individuals’ preferences for redistribution and explores the mechanisms through which this replication may fail. Our instruments come from the General Social Science (GSS), a major social‐science survey in the U.S. Preferences for redistribution are measured through both a standard question (asking whether the government should do everything possible to improve the standard of living of all poor Americans) and a health-tax willingness question. In addition, mechanism variables include fairness judgments, work-versus-luck beliefs, father’s education, and trust in others.
@@
Our investigation proceeds in two stages. First, we examine whether digital twins systematically differ from individuals in their preferences for redistribution. Second, we investigate whether any observed discrepancies can be explained by the twins’ failure to replicate responses related to socio‐economic heritage (father’s education), fairness judgments, beliefs about the relative role of effort versus luck in achieving success, and interpersonal trust. These have been shown to predict redistribution preferences [1].
@@
To understand why these mismatches occur, we examined additional constructs known to shape redistribution attitudes [1]. Twins poorly replicate socio-economic heritage—father’s education shows a low correlation ($r\approx.13$), a mean gap of over 1.5 years (t=21.55,p<.001t=21.55,\;p<.001), and a sharply reduced variance (F=12.18,p<.001F=12.18,\;p<.001). Fairness judgments ($r\approx.30$; accuracy $\approx.64$) and work-versus-luck beliefs ($r\approx.35$; accuracy $\approx.77$) similarly display moderate rank-order similarity but pronounced mean and variance discrepancies. Interpersonal trust also diverges, with twins overstating trust ($\text{mean}_{t}=2.22$ vs. $\text{mean}_{h}=1.73$; t=10.38,p<.001t=10.38,\;p<.001). Because these background factors and beliefs are important predictors of redistribution preferences, their imperfect replication by digital twins likely underlies the systematic errors in predicting absolute levels of support for redistribution.
@@
Political donations play a critical role in shaping electoral outcomes. They equip candidates with the financial resources needed to compete in political races (e.g., [19, 27]). They also function as signals of support. Political parties often use donation counts as thresholds for unlocking institutional backing and as a pre-requisite for qualifying for televised debates. In the present research, we investigate the reputational consequences of a specific form of donation behavior: contributing to both major political parties. We refer to individuals who donate to two parties as “promiscuous donors” and we examine how they are perceived by others. Understanding how promiscuous donors are perceived is important given the prevalence of this behavior. Analysis of 2024 Federal Election Commission data reveals that at least 113,187 Americans made contributions to both parties during the 2024 election cycle.
Ex ante, it is unclear how individuals will evaluate those who donate to both political parties. On one hand, promiscuous donors may be viewed favorably. Some scholars have argued that political polarization has reached unsustainable levels [24], prompting a public desire for greater bipartisanship and reconciliation. From this perspective, contributing to both parties could signal a conciliatory stance, potentially enhancing reputational standing.
On the other hand, several theoretical perspectives suggest that such behavior may incur reputational costs. Partisan animosity has intensified in recent years, with Democrats and Republicans expressing growing levels of dislike and distrust toward each other [47]. Furthermore, emerging research on receptiveness to opposing views [28] suggests that open-mindedness toward the other side can backfire socially [29]. If donors are seen as legitimizing or aligning with an ideologically objectionable outgroup, their behavior may be perceived as disloyal or suspect, thereby leading to reputational costs.
@@
We recruited 799 human respondents on Prolific and their respective twins. This study used a rating-based conjoint design [21]. Conjoint is a study design used to determine how people value different attributes (e.g., screen size, memory, color) that make up a product (e.g., a laptop). Participants in a conjoint study are shown a series of trials, each showcasing a different version of the target product. By varying different attribute levels of the product (e.g., a black vs. silver laptop) between trials and measuring how consumers respond to different combinations of these attributes, managers can quantify how much a given attribute (e.g., color) affects overall attitudes.
@@
"I currently donate to both Democrats and Republicans."
@@
How do people evaluate promiscuous donors (i.e., targets who donate to both parties)? Our results indicate that, compared to those who do not donate to politics at all, those who donate to both parties are perceived more negatively. This result is consistent with recent research on the reputational costs of receptiveness to opposing views [29]. Interestingly, the reputational costs of promiscuous donation are particularly salient when the reason behind donating to both parties is self-interest (e.g., to attract the best talent to one’s business or to make sure that one’s business can work with whichever party wins) and are somewhat weaker when the reason is more value-based (e.g., believing in bridging divides or that good ideas should be supported regardless of which party they come from).
@@
This study evaluates whether digital twins trained from large language models (LLMs) can accurately replicate individual responses on a newly developed psychometric instrument measuring Quantitative Intuition (QI) [18]. The QI scale assesses the respondent’s own analytical and intuitive mindsets through 38 self-report items in addition 16 items measuring individuals’ perceptions of their organization’s analytical and intuitive orientations.
@@
Companies commonly target their advertisements based on demographic characteristics, such as race or gender. Yet recent work has found that people often view the decision to target to be less fair than advertising broadly to the general population [53]. In this sub-study, we aim to replicate this pattern with human participants, and test whether their digital twins indicate similar fairness ratings.
@@
Modern machine learning and AI-based recommendation systems are data-driven, generating personalized recommendations from a user’s previous interactions. Often, platforms make the assumption that a user’s engagement with a particular piece of content is indicative of their utility for the content. However, a recent online behavioral experiment has shown that users may behave strategically and alter their engagement to influence the content they get recommended in the future [11]. For example, a user on a music platform may skip past a “guilty pleasure” song that they like, because they are worried that the recommendation algorithm will recommend too many similar songs later.
@@
For preference for user controls in recommendation systems, we ask “What controls, if any, would you like to have over their recommendation systems? Please check ALL that apply” with a menu of options. We count the number of options selected other than "No control: let the algorithm work automatically."

### v4

@@ Appendix H Centaur Experiments
    Self-persuasion (?). When prompted to consider the other side, do digital twins abandon their attitudes more readily than their human counterparts?
@@
In sum, across all three task types, we do not find evidence supporting the possibility that because twins obediently follow orders, they are more sensitive to survey instructions, such as ”consider the alternative point of view” or ”imagine…” If anything, we find that humans, on average, respond more to instructions in this survey than digital twins.
@@
Building on the seminal work of (?), this study assesses the ability of digital twins to replicate individuals’ preferences for redistribution and explores the mechanisms through which this replication may fail. Our instruments come from the General Social Science (GSS), a major social‐science survey in the U.S. Preferences for redistribution are measured through both a standard question (asking whether the government should do everything possible to improve the standard of living of all poor Americans) and a health-tax willingness question. In addition, mechanism variables include fairness judgments, work-versus-luck beliefs, father’s education, and trust in others.
@@
Our investigation proceeds in two stages. First, we examine whether digital twins systematically differ from individuals in their preferences for redistribution. Second, we investigate whether any observed discrepancies can be explained by the twins’ failure to replicate responses related to socio‐economic heritage (father’s education), fairness judgments, beliefs about the relative role of effort versus luck in achieving success, and interpersonal trust. These have been shown to predict redistribution preferences (?).
@@
To understand why these mismatches occur, we examined additional constructs known to shape redistribution attitudes (?). Twins poorly replicate socio-economic heritage—father’s education shows a low correlation ($r\approx.13$), a mean gap of over 1.5 years (t=21.55,p<.001t=21.55,\;p<.001), and a sharply reduced variance (F=12.18,p<.001F=12.18,\;p<.001). Fairness judgments ($r\approx.30$; accuracy $\approx.64$) and work-versus-luck beliefs ($r\approx.35$; accuracy $\approx.77$) similarly display moderate rank-order similarity but pronounced mean and variance discrepancies. Interpersonal trust also diverges, with twins overstating trust ($\text{mean}_{t}=2.22$ vs. $\text{mean}_{h}=1.73$; t=10.38,p<.001t=10.38,\;p<.001). Because these background factors and beliefs are important predictors of redistribution preferences, their imperfect replication by digital twins likely underlies the systematic errors in predicting absolute levels of support for redistribution.
@@
Political donations play a critical role in shaping electoral outcomes. They equip candidates with the financial resources needed to compete in political races (e.g., (?, ?)). They also function as signals of support. Political parties often use donation counts as thresholds for unlocking institutional backing and as a pre-requisite for qualifying for televised debates. In the present research, we investigate the reputational consequences of a specific form of donation behavior: contributing to both major political parties. We refer to individuals who donate to two parties as “promiscuous donors” and we examine how they are perceived by others. Understanding how promiscuous donors are perceived is important given the prevalence of this behavior. Analysis of 2024 Federal Election Commission data reveals that at least 113,187 Americans made contributions to both parties during the 2024 election cycle.
Ex ante, it is unclear how individuals will evaluate those who donate to both political parties. On one hand, promiscuous donors may be viewed favorably. Some scholars have argued that political polarization has reached unsustainable levels (?), prompting a public desire for greater bipartisanship and reconciliation. From this perspective, contributing to both parties could signal a conciliatory stance, potentially enhancing reputational standing.
On the other hand, several theoretical perspectives suggest that such behavior may incur reputational costs. Partisan animosity has intensified in recent years, with Democrats and Republicans expressing growing levels of dislike and distrust toward each other (?). Furthermore, emerging research on receptiveness to opposing views (?) suggests that open-mindedness toward the other side can backfire socially (?). If donors are seen as legitimizing or aligning with an ideologically objectionable outgroup, their behavior may be perceived as disloyal or suspect, thereby leading to reputational costs.
@@
We recruited 799 human respondents on Prolific and their respective twins. This study used a rating-based conjoint design (?). Conjoint is a study design used to determine how people value different attributes (e.g., screen size, memory, color) that make up a product (e.g., a laptop). Participants in a conjoint study are shown a series of trials, each showcasing a different version of the target product. By varying different attribute levels of the product (e.g., a black vs. silver laptop) between trials and measuring how consumers respond to different combinations of these attributes, managers can quantify how much a given attribute (e.g., color) affects overall attitudes.
@@
”I currently donate to both Democrats and Republicans.”
@@
How do people evaluate promiscuous donors (i.e., targets who donate to both parties)? Our results indicate that, compared to those who do not donate to politics at all, those who donate to both parties are perceived more negatively. This result is consistent with recent research on the reputational costs of receptiveness to opposing views (?). Interestingly, the reputational costs of promiscuous donation are particularly salient when the reason behind donating to both parties is self-interest (e.g., to attract the best talent to one’s business or to make sure that one’s business can work with whichever party wins) and are somewhat weaker when the reason is more value-based (e.g., believing in bridging divides or that good ideas should be supported regardless of which party they come from).
@@
This study evaluates whether digital twins trained from large language models (LLMs) can accurately replicate individual responses on a newly developed psychometric instrument measuring Quantitative Intuition (QI) (?). The QI scale assesses the respondent’s own analytical and intuitive mindsets through 38 self-report items in addition 16 items measuring individuals’ perceptions of their organization’s analytical and intuitive orientations.
@@
Companies commonly target their advertisements based on demographic characteristics, such as race or gender. Yet recent work has found that people often view the decision to target to be less fair than advertising broadly to the general population (?). In this sub-study, we aim to replicate this pattern with human participants, and test whether their digital twins indicate similar fairness ratings.
@@
Modern machine learning and AI-based recommendation systems are data-driven, generating personalized recommendations from a user’s previous interactions. Often, platforms make the assumption that a user’s engagement with a particular piece of content is indicative of their utility for the content. However, a recent online behavioral experiment has shown that users may behave strategically and alter their engagement to influence the content they get recommended in the future (?). For example, a user on a music platform may skip past a “guilty pleasure” song that they like, because they are worried that the recommendation algorithm will recommend too many similar songs later.
@@
For preference for user controls in recommendation systems, we ask “What controls, if any, would you like to have over their recommendation systems? Please check ALL that apply” with a menu of options. We count the number of options selected other than ”No control: let the algorithm work automatically.”